Beam4a,6b a2010 06,07oct11
- beam 4a has not been working since shortly after alfa was
reinstalled in aug11. It has strong oscillations.
- On 06oct11 (in the afternoon) the drain current was lowered
from 17 to 11 milliamps.
- The oscillations stopped and 4a started working.
- With the drain current = 11mamps, the Tsys was about 90 K.
- A 12 db amp was placed after the dewar to increase the power
level. Tsys decreased to about 55 K
- beam6b has intermittently been failing.
Data acquisition and processing:
The plots show the total power vs time for:
- a2010 took data on 06oct11 from 21:00 to about 3:00 (07oct11)
with beam4a and 6b both on.
- The total power was computed over each 100 Mhz band using the rms
by channel to remove rfi.
- 06oct11 (.ps) (.pdf)
- page 15 starts the evening observing for 06oct11 (after the
changes to 4a).
- The 12 sec peaks are the FAA radar which is very strong
- Both bm4a and bm6b are going into compression much more often
than the other beams.
- 4b is not compressing while 4a compresses with a 6 second
period (must be a second radar doing this).
- 6b compresses every 12 seconds while 6A does not compress.
- 07oct11 (.ps) (.pdf)
- 4a continues compressing.
- 6b starts to fail on page 1 around 00:13:34 am.This continues
the rest of the run.